Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 4,200 units, 2/3 completed 10,780
31 Direct materials, 75,600 units 151,200 161,980
31 Direct labor 43,210 205,190
31 Factory overhead 24,308 229,498
31 Goods finished, 76,500 units 221,710 7,788
31 Bal. ? units, 2/5 completed 7,788

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit $
2. Conversion cost per equivalent unit $
3. Cost of the beginning work in process completed during March $
4. Cost of units started and completed during March $
5. Cost of the ending work in process $

b. Assuming that the direct materials cost is the same for February and March, did the conversion cost per equivalent unit increase, decrease, or remain the same in March?
Increase
Answer #1

Ending work in process = 4200+75600-76500 = 3300 Units

Equivalent unit of material = (4200*0%+72300*100%+3300*100%) = 75600

Equivalent unit of conversion = (4200*1/3+72300*100%+3300*2/5) = 75020

a) Direct material Cost per equivalent unit = 151200/75600 = 2

b) Conversion Cost per equivalent unit = (43210+24308)/75020 = 0.90

c) Cost of beginning work in process Completed = (1400*.90+10780) = 12040

d) Cost of unit started and completed = 72300*2.9 = 209670

e) Cost of ending work in process = (3300*2+1320*.90) = 7788

2) Direct material Cost = 4200*2 = 8400

Conversion Cost = 10780-8400 = 2380

Conversion Cost per equivalent unit = 2380/2800 = 0.85

Increase by (0.90-.85) = 0.05